# Scientific Literature Hierarchy Construction and Organization

### v1

@@ 4.2 Top-down and Bottom-up Baselines
We explore special cases of Scychic: one that employs only the top-down approach and another that uses only the bottom-up approach. These variations allow us to assess the individual strengths and weaknesses of each strategy. We assess these specific cases in §5.4.
@@ 4.3 Pure LLM-based Baselines
Initializing a Seed Hierarchy: The first step involves creating a seed hierarchy, starting with the hierarchy of sciences from the Wikipedia page on branches of science. [^2] https://en.wikipedia.org/wiki/Branches_of_science We made several adjustments to this hierarchy, detailed in §D, where the resulting seed hierarchy is also included.
@@
4.4 The Computational Complexity of Each Approach
@@
An important bottleneck for building a hierarchy is the number of LLM calls. Let $C$ denote the total number of contributions (§3.2), $b$ the branching factor, and $L$ the maximum depth set to $L=O(\log_{b}C)$ for a near-balanced tree (§3.3). For our proposed algorithm, Scychic, both the bottom-up and top-down variants incur a total of $O\left(\frac{C}{b}+\frac{C}{b^{2}}+\dots+\frac{C}{b^{L}}\right)=O\left(C/b\right)$ LLM calls. This is because at each layer of the hierarchy, the number of nodes decreases geometrically by a factor of $b$. For the LLM-based baselines presented in §4.3: fLMSci (parallel) involves processing contributions in batches of size $l$ (in experiments we use $l=100$). Consequently, the total number of LLM calls is given by $O(C/l)$. In fLMSci (incremental), each of the $C$ items is inserted into the hierarchy by traversing from the root to a leaf. In a tree of depth $L=O(\log_{b}C)$, this traversal requires $O(\log_{b}C)$ steps per item, giving us a total number of LLM calls: $O(C\log_{b}C)$. In the experiments, we report the number of LLM calls for each approach where we see that the gap between $O(C\log_{b}C)$ and $O\left(C/b\right)$ is quite drastic (61K vs 322, in Table 1).
@@ 5 Experimental Setup and Results
We provide further context on our experimental setup: the collection of papers with diverse topics which we aim to place on a hierarchy (§5.1), and our evaluation setup (§5.2).
@@ 5.1 Collection of Science Papers
We compile a set of science papers to used in our study. Our primary analysis is conducted on a relatively small collection of roughly $2K$ papers, spanning domains such as computer science, neuroscience, biology, oceanography, and their interdisciplinary connections. We refer to this collection as SciPile. This smaller set allows for rapid iteration over the data, enabling us to test various algorithms and assess their scalability. These papers are used to validate our approach in generating high-quality hierarchies. To evaluate whether our methods can scale to larger collections and create a more comprehensive hierarchy, we curated an expanded dataset of $10K$ papers, referred to as SciPileLarge. Details on our collection and filtering is in §F.
@@ 5.2 Evaluation as Utilization
In an ideal world, we would compare the inferred hierarchy to a gold standard, but such data doesn’t exist, and scientific literature constantly evolves. Therefore, we need an evaluation framework independent of a fixed ground truth.
We propose assessing the quality and utility of generated hierarchies from a utilization perspective, focusing on their ability to aid navigation and identify relevant content. We use an LLM-based agent to locate target information within the hierarchy, performing tree traversal and measuring accuracy. Evaluation occurs at each level and across the entire hierarchy to determine overall success. A superior hierarchy enhances information discovery, capturing conceptual relationships among subtopics. While we evaluate based on its ability to locate information, this doesn’t restrict the hierarchy’s other uses.
Key considerations are (a) selecting suitable queries and (b) choosing the most accurate LM models. For (a), we sample a paper (title and abstract). We also considered using questions derived from the paper’s content to simulate typical information-seeking behavior. However, in our pilot studies, we found that results from both methods generally correlate, so we opted for the former. For (b), we use Qwen2.5-32b-instruct due to its strong instruction-following capabilities. In pilot experiments, Qwen demonstrated the highest consistency compared to other open-weight models and GPT-4 (see §B).
Starting at the top of the hierarchy, we present the evaluator (LLM) with our query and the clusters’ contributions (see prompt in Fig.2). The LLM must select the most appropriate cluster. We proceed to the next hierarchy level only if the model correctly identifies the cluster containing the target paper. At each subsequent level, we use the same query, presenting only the subclusters within the previously selected correct cluster. This process continues until the model successfully navigates to the level containing the target paper.
We report two metrics: Strict Accuracy (Strict-Acc), which measures how often the LLM correctly identifies the exact node, and Layer-1 Accuracy (L1-Acc), which measures how often the LLM correctly identifies the subtree containing the query.
Method Accuracy (%) LLM Cost Hierarchy Structure
Strict-Acc $\uparrow$ L1-Acc $\uparrow$
Avg. # of
Input Tokens $\downarrow$
# of
Calls $\downarrow$
Max
Depth
Avg
Depth
Avg Bran.
Factor
Max Bran.
Factor
# of
Items
Contributions type: Topic
Scychic 14.9 $\pm$ 2.7 65.7 $\pm$ 4.4 5017 322 3 3 40.9 128 11k
@@
18.0 $\pm$ 5.3 91.0 $\pm$ 4.0 4040 61K 14 7.7 3.6 704 10.4K
Table 1: Evaluation results of Scychic, fLMSci (parallel) and fLMSci(incremental) when using Topic as contribution type. “Bran.” stands for “Branching”. All methods show poor Strict-Acc ($\leq 18.0\%$), highlighting the challenging nature of the task. On one hand, fLMSci (inc) achieves the highest accuracy, showing the feasibility of building hierarchies by pure LLM-based methods. However, it incurs substantial computational costs, about 200$\times$ higher than other methods. In contrast, Scychic offers a balanced performance profile with competitive accuracy (14.9% Strict-Acc, 65.7% L1-Acc) while maintaining significantly lower computational requirements.
@@
Table 2: Evaluation results of Scychic and the corresponding baselines. Scychic maintains the highest accuracy and a relatively small variance across all contribution types, which proves the rationale behind our hybrid design. Among three contribution types, problem statement contribution type hierarchies have the highest Strict-Acc (51.1%) and near first L1-Acc(81.7%), indicating this contribution type contributes most for hierarchy construction.
5.3 Experiment Design
Following the proposed evaluation protocol, we conduct three sets of experiments to compare our method (Scychic, §4.1) with baselines (§4.2, 4.3). The hyperparameters of Scychic are in §G. The experiments are organized as follows: (1) We first compare all approaches on “topic” contributions, the simplest form of contributions. (2) Next, we assess the methods on more complex contributions (problem definition, results, and solutions). (3) Finally, the third set of experiments evaluates scalability on SciPileLarge. Due to the computational cost (number of API calls), we evaluate the LLM-based baselines only for the first set.
5.4 Empirical Findings
Tables 1, 2, and 3 present our experimental results. They also include LLM Cost, detailing the average number of input tokens and LLM calls, and Hierarchy Structure, which provides statistics on depth and branching factors (i.e., the number of children per node).
Scychic outperforms its special-cases baselines. From Table 1, we observe that Scychic significantly outperforms the top-down and bottom-up baselines in Level-1 accuracy while achieving comparable Strict-Acc. Similar outcomes are also see in Table 2 for contribution types. For instance, when on “solution” contributions, Scychic outperforms the top-down baseline by 2.9% in Strict-Acc and 3.1% in L1-Acc, demonstrating the effectiveness of our method. Moreover, these strong performances are achieved with a comparable number of tokens and LLM calls as the baselines, making Scychic computationally efficient.
LLM-based baselines are much costlier than Scychic. Although the fLMSci baseline outperform Scychic in both accuracies by a small margin, it requires an prohibitively large number of calls, rendering it too costly for large-scale scientific literature. Despite it’s good quality, fLMSci (incremental) will not scale.
Scychic scales to larger paper corpus. For our scalability experiments, we evaluated Scychic on our larger $10K$ paper dataset SciPileLarge, using the problem statement contribution type. Due to the significant increase (×$\times 5$) in corpus size, we implemented a 4-layer hierarchy instead of the 3-layer structure used previously. Notably, Scychic achieved even higher L1-Acc (86.5%) on SciPileLarge compared to our smaller dataset SciPile. This improvement likely stems from the enhanced quality of our expanded dataset, which has more strict filtering mechanisms. While the Strict-Acc showed a minor decrease compared to results on SciPile, it remained at a satisfactory level. Collectively, these results provide compelling evidence that our method scales successfully to substantially larger paper corpora.
Method Accuracy (%) LLM Cost Hierarchy Structure
Strict-Acc $\uparrow$ L1-Acc $\uparrow$
Avg. # of
Input Tokens $\downarrow$
# of
Calls $\downarrow$
Depth
(Max/Avg)
Avg Bran.
Factor
Max Bran.
Factor
# of
Items
@@ 5.4 Empirical Findings
26.2 $\pm$ 5.4 41.9 $\pm$ 4.0 5924 1572 4 8 26 $10K$
Table 3: Evaluation results of Scychic when building hierarchies on SciPileLarge (dataset with $10K$ papers). Compared to SciPile (the $2K$ dataset), our method shows a slight decrease in Strict-Acc, while even achieves higher L1-Acc. This fully demonstrates the feasibility of our method on the scaling dataset.
@@ 6 Discussion and Conclusion
Future applications: Our work paves the way for several promising research directions. One potential application is to leverage the constructed hierarchies as tools for exploratory analysis in various scientific fields. These hierarchies can help academic institutions and funding agencies identify emerging trends and under-explored areas, and they can be refined for domain-specific analyses that capture the nuances of individual disciplines. This approach not only enriches our understanding of scientific progress but also offers a new perspective on organizing the vast landscape of scholarly work.
Limitations: Although we evaluated our pipeline on $10K$ papers, this is still far from the true scale of scientific literature. We hope future work will enhance our approach to handle more realistic scales. Additionally, while our evaluation framework shows potential for efficient information discovery, it may have its own weaknesses and biases. Integrating human verification into the assessment process could help ensure the quality and reliability of the inferred hierarchies.
Conclusions: In summary, we have motivated Science Hierarchography, large-scale hierarchical summarization of scientific literature, offering a fresh perspective on how research efforts are organized and distributed. Our method, Scychic, effectively combines LLMs with faster procedures to balance quality and scalability. Looking ahead, we hope our work offers researchers a more intuitive way to navigate the complex landscape of scholarly and enhances the allocation of resources in academia.
@@
Acknowledgments
This work is supported by ONR grant (N0001424-1-2089) and Defense Advance Research Projects Agency (DARPA) under Contract No. HR001125C0304. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of DARPA. We sincerely thank Jiefu Ou and the broader JHU CLSP community for discussions and inspiration.
@@ Appendix A Additional Related Work
We include additional related work here, because of the space limitation in the main text.
@@ C.1 Prompt for Extracting Problem/Solution/Result Contributions
We use the prompt below to extract contributions from the paper’s title and abstract. After finishing the extraction, the three contributions will be saved into the original json file.
@@ C.3 Examples for Problem/Solution/Result/Topic contributions extracted from papers
Abstract: Due to the lack of full rotational symmetry in condensed matter physics, solids exhibit new excitations beyond Dirac and Weyl fermions, of which the sixfold excitations have attracted considerable interest owing to the presence of maximum degeneracy in bosonic systems. Here, we propose that a single linear dispersive sixfold excitation can be found in the electride Li12 Mg3 Si4 and its derivatives. The sixfold excitation is formed by the floating bands of elementary band representation A@12a originating from the excess electrons centered at the vacancies (i.e., the 12a Wyckoff sites). There exists a unique topological bulk-surface-edge correspondence for the spinless sixfold excitation, resulting in trivial surface “Fermi arcs” but topological hinge arcs. All gapped kz slices belong to a two-dimensional higher-order topological insulating phase, which is protected by a combined symmetry T S4z and characterized by a quantized fractional corner charge Qcorner = 3||||e||||/4. Consequently, the hinge arcs are obtained in the hinge spectra of the S4z-symmetric rod structure. The state with a single sixfold excitation, stabilized by both nonsymmorphic crystalline symmetries and time-reversal symmetry, is located at the phase boundary and can be driven into various topologically distinct phases by explicit breaking of symmetries, making these electrides promising platforms for the systematic studies of different topological phases.
[…]
@@
{ "overarching_problem_domain": "Condensed matter physics", "challenges/difficulties": "Lack of full rotational symmetry in solids leading to new excitations beyond Dirac and Weyl fermions", "research_question/goal": "Investigate sixfold excitations in electrides" } { "overarching_solution_domain": "Electrides and topological phases", "solution_approach": "Propose that a single linear dispersive sixfold excitation can be found in the electride Li12Mg3Si4 and its derivatives", "novelty_of_the_solution": "Unique topological bulk-surface-edge correspondence for the spinless sixfold excitation" } { "findings/results": "The sixfold excitation is formed by floating bands of elementary band representation A@12a. All gapped kz slices belong to two-dimensional higher-order topological insulating phase, characterized by a quantized fractional corner charge Qcorner = 3||||e||||/4. Hinge arcs are obtained in the hinge spectra of the S4z-symmetric rod structure.", "potential_impact_of_the_results": "Electrides are promising platforms for systematic studies of different topological phases." }
[…]
@@
[…]

### v2

@@ 4.2 Top-down and Bottom-up Baselines
We examine two special cases of Scychic: one using only a top-down strategy and the other solely with a bottom-up approach. These variants help isolate and evaluate the strengths and limitations of each method. Results are discussed in §5.4.
@@ 4.3 Pure LLM-based Baselines
Initializing a Seed Hierarchy: The first step involves creating a seed hierarchy, starting with the hierarchy of sciences from the Wikipedia page on branches of science.222en.wikipedia.org/wiki/Branches_of_science We made several adjustments to this hierarchy, detailed in §D, where the resulting seed hierarchy is also included.
@@
4.4 Computational Complexity of Approaches
A major scalability bottleneck in hierarchy construction is the number of LLM calls. Let $C$ be the number of contributions (§3.2), $b$ the branching factor, and $L=O(\log_{b}C)$ the maximum depth for a near-balanced tree (§3.3). Our proposed algorithm, Scychic, requires $O\left(C/b\right)$ LLM calls for both its top-down and bottom-up variants. Among the LLM-based baselines discussed in §4.3, fLMSci (parallel) makes $O(C/l)$ calls (with $l$ as batch size), offering lower complexity but at the cost of reduced quality. In contrast, fLMSci (incremental) achieves higher accuracy but requires $O(C\log_{b}C)$ LLM calls due to root-to-leaf traversals during insertion. Empirically, the difference in LLM usage is significant: in our 2K-paper setup, fLMSci (incremental) makes 61K calls compared to just 322 for Scychic (subsection 5.3).
@@ 4.4 Computational Complexity of Approaches
Table 2: Computational complexity of hierarchy construction methods measured by LLM calls, with $C$ = contributions, $b$ = branching factor, and $l$ = batch size.
@@ 5 Experimental Setup and Results
We describe our experimental setup, including the diverse paper collection used for our experiments (§5.1) and the evaluation framework (§5.2).
@@ 5.1 Collection of Science Papers
We compile a collection of scientific papers spanning domains such as computer science, neuroscience, biology, oceanography, and their interdisciplinary intersections. Our initial analysis focuses on a smaller set of approximately $2K$ papers (referred to as SciPile), allowing for rapid iteration over design choices and assessment of scalability. We then extend our analysis to a larger collection of $10K$ papers, referred to as SciPileLarge. Details on data collection and filtering are provided in §F.
@@ 5.2 Evaluation as Utilization
Ideally, hierarchy quality would be evaluated against a gold standard—but no such reference exists, and scientific literature continually evolves. As a result, we adopt an evaluation framework based on utilization, independent of fixed ground truth.
We assess hierarchy quality by measuring how well it supports navigation and content discovery. Specifically, we use an LLM-based agent to locate target papers via tree traversal, tracking accuracy at each level and across the full hierarchy. A stronger hierarchy should better capture conceptual relationships and improve information-seeking efficiency. While our evaluation focuses on retrieval, the hierarchy’s utility extends beyond that.
Two key design choices guide our evaluation: (a) selecting appropriate queries, and (b) choosing a reliable LLM. For (a), we sample paper titles and abstracts as queries. Although we considered generating language questions from papers, pilot studies showed both approaches yield similar results, so we use the simpler method. For (b), we use Qwen2.5-32b-instruct, which demonstrated performance closest to GPT-4 among open models in our pilot evaluations (§B).
Starting at the top of the hierarchy, we provide the evaluator (LLM) with the query and the list of cluster descriptions (prompt in Fig. 2). The model must select the cluster it deems most relevant. We move to the next level only if the chosen cluster contains the target paper. At each subsequent level, the same query is used, but only subclusters of the previously selected (and correct) cluster are shown. This continues until the model reaches the level containing the target paper. We report two metrics: Strict Accuracy (Strict-Acc), which captures how often the model selects the exact correct node, and Layer-1 Accuracy (L1-Acc), which measures how often it selects the correct top-level subtree.
5.3 Experiment Design
We conduct few sets of experiments to evaluate our method (Scychic, §4.1) against the baselines (§4.2, 4.3) following the proposed evaluation protocol. Hyperparameter details for Scychic are provided in §G. The experiments are structured as follows: (1) We begin by comparing all methods on “topic” contributions, the simplest contribution type (Table 5.3). Due to the high computational cost, LLM-based baselines are evaluated only in this setting. (2) We then evaluate performance on more complex contributions (problem, solution, and results) using both SciPile and SciPileLarge to test scalability (Table 4). The results tables also report LLM Cost (average number of input tokens and LLM calls) and Hierarchy Structure (depth and branching factor per node).
Method Strict-Acc (%) $\uparrow$ L1-Acc (%) $\uparrow$ # of Calls $\downarrow$
Topic contributions
Scychic 14.9 $\pm$ 2.7 65.7 $\pm$ 4.4 322
@@ 5.3 Experiment Design
18.0 $\pm$ 5.3 91.0 $\pm$ 4.0 61K
[…]
@@
Dataset: SciPileLarge (10K papers)
@@
26.2 $\pm$ 5.4 41.9 $\pm$ 4.0 5924 26
Contributions type: Solution Statement
Scychic 24.7 $\pm$ 4.8 65.8 $\pm$ 2.5 7653 28
[…]
Top-down
22.4 $\pm$ 3.5 52.3 $\pm$ 3.0 4032 1572 4 8 26
[…]
Bottom-up
23.9 $\pm$ 3.3 51.3 $\pm$ 3.1 6150 28
Contributions type: Results Statement
Scychic 27.6 $\pm$ 4.6 69.8 $\pm$ 2.1 6457 30
[…]
Top-down
19.7 $\pm$ 4.0 54.0 $\pm$ 3.3 5380 1572 4 8 30
[…]
Bottom-up
23.6 $\pm$ 2.7 55.2 $\pm$ 2.9 4731 28
Table 4: Evaluation results of Scychic and the corresponding baselines on both the 2K (SciPile) and 10K (SciPileLarge) datasets. Scychic maintains high accuracy and a relatively small variance, proving the rationale behind our hybrid design. When scaling from 2K to 10K papers, our method shows a slight decrease in Strict-Acc but maintains strong L1-Acc, demonstrating its feasibility on larger datasets. Across both scales, the problem statement contribution type consistently yields the most accurate hierarchies, indicating this contribution type contributes most for hierarchy construction.
5.4 Empirical Findings
Scychic outperforms its special-case baselines. As shown in Table 5.3, Scychic achieves higher Level-1 accuracy than the top-down and bottom-up baselines, while maintaining comparable Strict-Acc. Similar trends hold across other contribution types in Table 4. For example, on “solution” contributions, Scychic exceeds the top-down baseline by 2.9% in Strict-Acc and 3.1% in L1-Acc, highlighting its effectiveness. Notably, these gains are achieved with a similar number of tokens and LLM calls, underscoring Scychic’s compute efficiency.
LLM-based baselines can be far more expensive than Scychic. While fLMSci slightly outperforms Scychic in accuracy, it does so at the cost of a massive increase in LLM calls—making it impractical for large-scale use. As a result, despite its strong performance, fLMSci (incremental) simply doesn’t scale.
Scychic scales to larger paper corpus. For our scalability experiments, we evaluated Scychic on our larger $10K$ paper dataset SciPileLarge, using the problem statement contribution type. Due to the significant increase ($\times 5$) in corpus size, we implemented a 4-layer hierarchy instead of the 3-layer structure used previously. Notably, Scychic achieved even higher L1-Acc (86.5%) on SciPileLarge compared to our smaller dataset SciPile. This improvement likely stems from the enhanced quality of our expanded dataset, which has more strict filtering mechanisms. While the Strict-Acc showed a minor decrease compared to results on SciPile, it remained at a satisfactory level. Collectively, these results provide compelling evidence that our method scales successfully to substantially larger paper corpora.
5.5 Additional Analyses
We briefly cover additional analyses that were omitted from the main text due to space constraints.
Detailed prompts significantly improve hierarchy quality. To demonstrate this, we compare two prompt types. The first is a "detailed" prompt—carefully curated with comprehensive instructions and reminders—which we use for all main experiments in this paper. The second is a "simplified" prompt containing only the core task description. The results confirm that the detailed prompt consistently and substantially outperforms the simplified version across all scenarios. More detailed results are in §H.3.
Embedding quality varies significantly across models. For the embedder mentioned in §4.1. We evaluated three models—Qwen’s gte-Qwen2-7B-instruct (Li et al., 2023), OpenAI’s text-embedding-3-large, and text-embedding-ada-002. The first two models perform similarly, whereas text-embedding-ada-002 produces markedly weaker results. We select gte-Qwen2-7B-instruct for its strong balance of performance and its practical value as an open-weight model for reproducible research. The experimental results are in §H.2.
5.6 Sample Visualization of the Hierarchy
The reader might be curious to see the resulting hierarchies. In §I we show a slice of the final hierarchy generated by Scychic on the SciPileLarge  dataset. The original hierarchy has 4 levels, use papers’ problem contribution. Due to space constraints, this slice shows only two levels of clusters above the individual papers.
@@ 6 Discussion and Conclusion
Future applications:
Our work opens several promising directions for future research. One key opportunity is to use the constructed hierarchies as tools for exploratory analysis across scientific domains. They can aid academic institutions and funding bodies in identifying emerging trends and underexplored areas, and can be adapted for domain-specific analyses that capture the unique structure of individual fields. This approach not only deepens our understanding of scientific progress but also provides a new lens for organizing the vast and growing body of scholarly work.
Conclusions:
We introduced Science Hierarchography, a framework for large-scale hierarchical summarization of scientific literature, offering a new lens on how research efforts are distributed. Our method, Scychic, combines LLMs with efficient algorithms to strike a balance between quality and scalability. Looking forward, we aim for this work to help researchers navigate the scientific landscape more intuitively and support more informed resource allocation in academia.
Limitations
Although we evaluated our pipeline on $10K$ papers, this is still far from the true scale of scientific literature. We hope future work will enhance our approach to handle more realistic scales. Additionally, while our evaluation framework shows potential for efficient information discovery, it may have its own weaknesses and biases. Integrating human verification into the assessment process could help ensure the quality and reliability of the inferred hierarchies.
@@ Appendix A Additional Related Work
We include additional related work here because of the space limitation in the main text.
@@ C.1 Prompt for Extracting Problem/Solution/Result Contributions
We use the prompt below to extract contributions from the paper’s title and abstract. After finishing the extraction, the three contributions will be saved into the original json file. Please see §3.2 for more information.
@@ C.3 Examples for Problem/Solution/Result/Topic contributions extracted from papers
Abstract: Due to the lack of full rotational symmetry in condensed matter physics, solids exhibit new excitations beyond Dirac and Weyl fermions, of which the sixfold excitations have attracted considerable interest owing to the presence of maximum degeneracy in bosonic systems. Here, we propose that a single linear dispersive sixfold excitation can be found in the electride Li12 Mg3 Si4 and its derivatives. The sixfold excitation is formed by the floating bands of elementary band representation A@12a originating from the excess electrons centered at the vacancies (i.e., the 12a Wyckoff sites). There exists a unique topological bulk-surface-edge correspondence for the spinless sixfold excitation, resulting in trivial surface “Fermi arcs” but topological hinge arcs. All gapped kz slices belong to a two-dimensional higher-order topological insulating phase, which is protected by a combined symmetry T S4z and characterized by a quantized fractional corner charge Qcorner = 3||e||/4. Consequently, the hinge arcs are obtained in the hinge spectra of the S4z-symmetric rod structure. The state with a single sixfold excitation, stabilized by both nonsymmorphic crystalline symmetries and time-reversal symmetry, is located at the phase boundary and can be driven into various topologically distinct phases by explicit breaking of symmetries, making these electrides promising platforms for the systematic studies of different topological phases.
@@
{ "overarching_problem_domain": "Condensed matter physics", "challenges/difficulties": "Lack of full rotational symmetry in solids leading to new excitations beyond Dirac and Weyl fermions", "research_question/goal": "Investigate sixfold excitations in electrides" } { "overarching_solution_domain": "Electrides and topological phases", "solution_approach": "Propose that a single linear dispersive sixfold excitation can be found in the electride Li12Mg3Si4 and its derivatives", "novelty_of_the_solution": "Unique topological bulk-surface-edge correspondence for the spinless sixfold excitation" } { "findings/results": "The sixfold excitation is formed by floating bands of elementary band representation A@12a. All gapped kz slices belong to two-dimensional higher-order topological insulating phase, characterized by a quantized fractional corner charge Qcorner = 3||e||/4. Hinge arcs are obtained in the hinge spectra of the S4z-symmetric rod structure.", "potential_impact_of_the_results": "Electrides are promising platforms for systematic studies of different topological phases." }
Contribution - Topic: ’Electrides’, ’Electrides in Condensed Matter Physics’, ’Higher-Order Topological Insulators’, ’Nonsymmorphic Symmetries’, ’Sixfold Excitation in Solids’, ’Sixfold Excitations’, ’Symmetry Breaking in Topological Materials’, ’Topological Bulk-Surface-Edge Correspondence’, ’Topological Phase Transitions’, ’Topological Phases in Condensed Matter Physics’, ’Topological Properties’
@@
Contribution - Topic: ’Convergence Testing in Computational Simulations’, ’Density Functional Theory (DFT) Accuracy’, ’High-Throughput Computational Methods’, ’Hubbard U Corrections’, ’Tin Phase Transitions’